OpenStudy (anonymous):

Can someone help me with stoichiometry?? I only have tonight to complete it and I need help Dx

OpenStudy (aaronq):

Stoichiometry When performing stoichiometric calculations the values to be used \(\sf \color{red}{must~be~in~moles}\)*. This is because the Molar Masses (\(\sf M_{m}\) of two elements are not the same. This is to say that 1 gram of Hydrogen atoms is not the same number of atoms as 1 gram of Nitrogen atoms. We need to be working with numbers of particles (e.g. atoms, molecules, formula units) and using moles is the way this is done. *A shortcut can be taken when working with \(\sf \color{blue}{Ideal~Gases}\) as moles are \(\propto\) volume. And, so volume can be regarded as moles, where 22.414 L = 1 mol. General Scheme: \(\sf \large 1.\)First write and balance an equation for the process described. \(\sf \large 2.\)Next, use the stoichiometric coefficients to find moles produced. \(\sf \large 3.\) Set up a ratio using the species of interest, like so (for a general reaction): \(\sf \large \color{red}{a}A + \color{blue}{b}B \rightleftharpoons \color{green}{c}C\) here upper case letters (A,B,C) are the chemical species, and lower case letters (a,b,c) are the coefficients, \(\sf \dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) You can use any pair from here, depending on what the question is asking for. For example: if you have 2 moles of B, how many moles of C can you produce? Now rearrange algebraically, \(\sf\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) \(\sf \large 4.\) Solve ———————————————